### v7

@@ 2 Related Works
2.1 Applications of CNN-based Model in Power Line Defect Detection
Yang et al. Yang et al. (2023) proposed DRA-Net, a dual-branch residual attention CNN-based model for pixel-wise power line detection, incorporating residual and attention modules to enhance feature representation, which significantly improved detection accuracy; however, its performance still degrades in extremely complex backgrounds due to the model’s limited ability to generalize across highly cluttered scenes.Liu et al.Liu et al. (2021)proposed the Box-Point Detector, a CNN-based insulator fault diagnosis method that integrates object detection and keypoint localization to improve the precision of defect identification in aerial imagery, achieving high detection accuracy and robustness; however, its reliance on manual feature labeling and limited evaluation on diverse environmental conditions may constrain its scalability and generalization in large-scale power line inspections. Yu et al.Yu et al. (2023) developed RcpVGG-YOLOv5, a lightweight YOLOv5-based insulator defect detection algorithm that integrates noise reduction and objects detection to enhance robustness in complex aerial backgrounds, achieving strong engineering applicability with reduced false detections; however, its lack of dedicated small objects detection strategies and the limited dataset size constrained accuracy to below 80%, indicating room for improvement in generalization and precision. Zhang et al.Zhang et al. (2025) introduced FAFNet, a CNN-based model incorporating feature alignment and multi-scale feature fusion to detect overhead transmission line components, which significantly enhanced detection accuracy and robustness; however, the method still faces limitations in handling small or heavily occluded objects, primarily due to the insufficient feature representation capability under extreme visual conditions. Zhang et al.Zhang et al. (2021)developed InsuDet, a CNN-based fault detection model for insulators in overhead transmission lines, integrating feature enhancement and region proposal improvements to achieve higher detection accuracy; nevertheless, its performance deteriorates under severe occlusion and complex backgrounds, mainly due to the insufficient robustness of feature extraction in highly cluttered environments. Tunio et al.Tunio et al. (2024)proposed a CNN-based model for fault detection and classification in overhead transmission lines, integrating comprehensive feature extraction and optimization strategies to enhance accuracy, which achieved superior classification performance; however, the model still struggles with generalization under highly variable fault conditions, primarily due to the limited adaptability of extracted features to unseen scenarios.
@@ 2.2 Applications of Transformer-Based Model in Power Line Defect Detection
3 Materials and Methods
@@ 3.2 Edge Enhanced ResNet
Formally, let the three convolutional branches be denoted as , corresponding to CDC, HDC, VDC and the conventional convolution, respectively. During training, their weights and biases are linearly combined:
@@ 3.4 Cross-Stage Dual-Domain Multi-Scale Attention Module
3.5 Focaler-Wise-SIoU Loss Function
In transmission line defect detection, most objects are extremely small and exhibit subtle spatial variations, which makes precise bounding box regression essential for reliable detection. Intersection-over-Union (IoU)-based losses are widely used in modern detectors for their geometric interpretability and optimization effectiveness, yet they still suffer from fundamental limitations. The canonical IoU loss, which measures only the overlap ratio between predicted and ground-truth boxes, fails to provide informative gradients when the boxes do not intersect, hindering optimization in challenging cases. Extensions such as GIoU (Rezatofighi et al., 2019), DIoU (Zheng et al., 2020), and CIoU (Zheng et al., 2022) address this issue partially by incorporating geometric constraints (e.g., enclosure area, center distance, aspect ratio), but they remain limited in jointly modeling spatial alignment and shape consistency. More advanced designs, such as SIoU (Gevorgyan, 2022), integrate angular, distance, and shape factors, yet they typically treat all samples equally and fail to account for difficulty imbalance, causing easy samples to dominate the optimization while hard ones—particularly small, low-SNR objects—are underemphasized. Moreover, conventional IoU-based losses lack discriminative power in high-IoU regimes, where sub-pixel localization precision is critical, resulting in slow convergence and degraded performance.
To overcome these challenges, we propose Focaler-Wise-SIoU Loss, a unified regression objective that integrates normalized IoU scaling, comprehensive geometric penalties, and a non-monotonic focal modulation mechanism. The proposed formulation enhances gradient sensitivity in high-IoU regions, improves geometric modeling fidelity, and adaptively reweights samples based on their difficulty, thereby accelerating convergence and significantly improving detection accuracy for small, hard-to-localize objects.
Formally, given a predicted bounding box and a ground-truth box , parameterized by their center coordinates and width–height pairs, we first convert them into the corner representation . The intersection and union areas are computed to yield the standard IoU metric:
To improve discriminability in high-IoU regions, we apply a linear normalization known as Focal-IoU scaling, mapping into the interval as:
where and . The base IoU penalty is then defined as:
On top of this formulation, we integrate the Scylla-IoU (SIoU) penalty, which accounts for angular, distance, and shape discrepancies. The angular component is derived from the displacement vector between box centers :
This term modulates the distance penalty as follows:
where denote the width and height of the smallest enclosing box. To further capture shape consistency, we define the shape penalty as:
where . The resulting SIoU-based core loss is expressed as:
To adaptively account for sample difficulty, particularly in datasets dominated by small and hard-to-localize objects, we introduce a Wise Modulation mechanism. A relative difficulty factor is defined as:
where is an exponential moving average updated with a momentum of . Unlike conventional monotonic modulation strategies, which monotonically emphasize harder samples and risk overfitting to outliers, we employ a non-monotonic focal modulation function:
where and are modulation hyperparameters. This formulation selectively amplifies moderately difficult samples while suppressing trivial or excessively noisy cases, contributing to more stable and balanced training dynamics.
Finally, the complete Focaler-Wise-SIoU Loss is defined as:
This unified formulation preserves the geometric interpretability and simplicity of traditional IoU-based losses while significantly enhancing discriminative capacity in high-IoU regimes, adaptively balancing sample contributions, and stabilizing the optimization process. As a result, the proposed Focaler-Wise-SIoU Loss demonstrates superior robustness and precision in transmission line defect detection, where most objects are small, difficult to localize, and require highly accurate bounding box regression.
4 Evaluation and results analysis
4.1 CSG-ADCD
@@
The CSG-ADCD (China Southern Grid Aerial Defective Component Dataset) is a large-scale, high-resolution UAV-based image dataset specifically developed for the detection of defects in power transmission line components, including insulators, tie wires, and poles. Constructed under authentic inspection scenarios by China Southern Power Grid, the dataset comprises 10,000 aerial images. Each image was manually annotated by a single domain expert to ensure both consistency and high-quality labeling.
@@
In summary, the CSG-ADCD dataset demonstrates strong representativeness by simultaneously addressing two critical aspects. On the one hand, it captures a comprehensive range of component defects and status indicators that are directly relevant to transmission safety, while also reflecting the complexity of real-world inspection conditions such as occlusion, cluttered backgrounds, and varying illumination. On the other hand, its pronounced dominance of small-scale objects—over 94% of all annotated instances—together with the presence of subtle and fine-grained anomalies, makes it a highly challenging and domain-specific benchmark. In this regard, CSG-ADCD not only advances research in defect detection for power systems but also provides a valuable testbed for generic small-object detection in aerial remote sensing, offering greater domain relevance than existing datasets such as VisDrone(Zhu et al., 2022).
@@ 4.2 Model Training and Evaluation Metrics
We implement the proposed model on the PyTorch framework and build it upon the Ultralytics RT-DETR backbone initialized with pretrained weights. We conduct training on a workstation equipped with an NVIDIA GeForce RTX 3090 GPU (24 GB) running Ubuntu 20.04, Python 3.10.14, PyTorch 2.3.1, and CUDA 12.1. We train the model for 100 epochs with a batch size of 4 and an input resolution of , using the AdamW optimizer with an initial learning rate of , a final learning rate factor , momentum of 0.9, and weight decay of . We apply a warm-up of 2000 iterations with momentum set to 0.8 and bias learning rate set to 0.1. For evaluation, we use the validation split following the COCO protocol with an IoU threshold of 0.7 and a maximum of 300 detections per image.
@@ 4.4 Ablation Study of the Proposed Method
To analyze the impact of each innovative component in CSG-ADCD, we conducted ablation experiments by systematically including and excluding four key modules: EER (Edge-Enhanced ResNet), SPD (Stride-free Space-to-Depth), CSDMAM (Cross-Stage Dual-Domain Multi-Scale Attention Module), and FWS Loss (Focaler-Wise-SIoU Loss). The results, shown in Table 2, demonstrate the contribution of each module to the model’s overall performance across various evaluation metrics.
@@
Edge Enhanced ResNet (EER) As shown in the table, enabling Edge Enhanced ResNet significantly improves the model’s precision and recall. Precision increases from 0.369 to 0.444, while recall increases from 0.177 to 0.203. This suggests that Edge Enhanced ResNet effectively refines error estimation, leading to better model accuracy, particularly in complex scenarios. Additionally, mAP50 improves from 0.163 to 0.189, further reflecting the positive impact of this module on object detection. Given that transmission line defect detection typically involves small, often faint features in images, EER plays a crucial role in enhancing the model’s sensitivity to these subtle defects.
Space-to-depth Convolution (SPD) Incorporating Space-to-depth Convolution leads to further performance improvements. Precision increases from 0.444 to 0.486, while recall remains relatively consistent at 0.195. The mAP50 also improves from 0.189 to 0.190, which shows that Space-to-depth Convolution is effective in fine-tuning object localization and enhancing precision without a significant trade-off in recall. Given the small scale of defects in transmission line imagery, SPD effectively helps capture more fine-grained spatial information, which improves the localization of small defects in the images.
Cross-Stage Dual-Domain Multi-Scale Attention Module (CSDMAM) The introduction of CSDMAM demonstrates its effectiveness in enhancing the model’s capability to capture small-scale defects under complex backgrounds. Although an intermediate configuration without the loss function shows minor fluctuations (e.g., precision 0.425 vs. 0.417, recall 0.190 vs. 0.208), the overall performance trend confirms the complementary role of CSDMAM when integrated with other modules. Notably, when CSDMAM is combined with both EER and SPD, the improves markedly from 0.083 to 0.119, highlighting its strength in boosting sensitivity to small objects. This suggests that CSDMAM effectively suppresses background interference while enhancing multi-scale feature representation, which is crucial for distinguishing subtle defects from visually similar structures in transmission line imagery. The observed gains in validate the necessity of CSDMAM for robust small-target detection, particularly when used synergistically with the proposed loss function.
Focaler-Wise-SIoU Loss (FWS Loss) The inclusion of Focaler-Wise-SIoU Loss leads to significant improvements across all metrics. Precision increases from 0.458 to 0.500, and recall increases from 0.166 to 0.204. mAP50 improves from 0.165 to 0.207, confirming that this loss function enhances model robustness, especially in challenging detection tasks. The APs also shows a marked improvement, increasing from 0.071 to 0.119. This improvement is especially important for the transmission line defect detection task, where small defects make up the majority of the dataset. The FWS Loss function is particularly effective in boosting the model’s sensitivity to small object detection without compromising overall performance.

### v8

@@ 2.1 Applications of Transformer-Based Model in Power Line Defect Detection
2.2 Joint Frequency–Spatial Representations for Small Object Detection
A growing body of work argues that purely spatial attention is insufficient for detecting small, weak, and cluttered targets, and instead advocates joint frequency–spatial representations that explicitly preserve high-frequency cues while suppressing background aliasing and downsampling loss. Representative approaches inject wavelet- or spectrum-domain operators into DETR/YOLO backbones: for instance, replacing standard operators with wavelet-domain anti-aliasing and convolution improves fine-structure perception and multi-scale representation under real-time constraints Shao et al. (2025). In infrared small-target detection, Haar wavelet decompositions have been used to derive high-/low-frequency masks that selectively retain fine details and mitigate over-smoothing, achieving state-of-the-art performance and illustrating an “frequency-guides-spatial” paradigm Feng et al. (2025); subsequent designs integrate frequency–spatial convolutions and cross-domain attention to highlight local high-frequency details while using low frequencies to complete global structure, yielding consistent gains with lightweight models Xu et al. (2025). From the feature fusion perspective, frequency-gated attention strengthens high-frequency details that are easily drowned out by deep downsampling, and combining it with grouped spatial–channel transformers produces robust small-target descriptors Li et al. (2025); similarly, explicitly separating and re-fusing spatial and frequency channels in hierarchical transformers alleviates the loss of detail caused by aggressive subsampling and semantics-dominant deep layers Wang et al. (2025). Importantly, these studies also caution that frequency interaction is not “the more the better”: without proper masking or dynamics modeling, it may amplify boundary noise; coupling frequency-interaction attention with stabilizing mechanisms has been shown to suppress irrelevant high-frequency artifacts and emphasize targets nearly indistinguishable from background Chen and Ye (2024). Beyond wavelets, discrete cosine transform (DCT) features, aligned and fused with spatial cues, have improved AP for small objects in remote sensing, indicating that cosine-spectrum bases offer complementary discriminative power for tiny structures Luo et al. (2024). Earlier frequency/multi-resolution methods (e.g., NSCT/Contourlet families) have also informed background estimation and low-frequency isolation: isolating low-frequency components reduces sensor and clutter noise while high-frequency components recover target details, providing a principled route to boost the small-target signal-to-noise ratio Li et al. (2024); Liu et al. (2018). In UAV scenarios emphasizing on-board efficiency, integrating wavelet convolutions and lightweight attention into real-time DETR variants demonstrates stable gains on diminutive, high-altitude objects and underscores the efficacy of “frequency-aware convolutions + transformers” under embedded constraints Guo et al. (2025).
In summary, the literature converges on three lessons. First, frequency decomposition should be paired with spatial alignment or masking to suppress noisy high-frequency components Feng et al. (2025); Chen and Ye (2024). Second, coupling with multi-scale or large-receptive-field modules is crucial to balance local details and global context Shao et al. (2025); Xu et al. (2025). Finally, the practical impact depends on lightweight implementations suitable for UAV platformsGuo et al. (2025).
3 Methods
@@ 3.2 Edge Enhanced ResNet
Formally, let the four convolutional branches be denoted as , corresponding to CDC, HDC, VDC, and the Vanilla convolution, respectively. During training, their weights and biases are linearly aggregated to obtain an equivalent kernel and bias:
@@ 3.4 Cross-Stage Dual-Domain Multi-Scale Attention Module
After an extensive literature review, we find that coupling frequency decomposition with attention is particularly effective for tiny, edge-dominated defects in UAV imagery—preserving high-frequency details and suppressing low-frequency clutter that misleads purely spatial attention. Therefore, we propose CSDMAM, which injects learnable frequency gating directly into the attention formation while retaining anisotropic, large-kernel spatial paths tailored to line-shaped defects.
@@
3.5 Focaler–Wise–SIoU Loss
In transmission-line defect detection, targets are extremely small and exhibit subtle spatial variations, so precise bounding-box regression is pivotal for reliable recognition. IoU-based losses are attractive for their geometric interpretability and optimization efficiency, yet the canonical IoU measures only overlap and provides no informative gradients when the boxes are disjoint, which stalls learning on difficult examples. Subsequent variants—GIoU Rezatofighi et al. (2019), DIoU Zheng et al. (2020), and CIoU Zheng et al. (2022)—mitigate this to some extent by adding enclosure, center-distance, and aspect-ratio terms. Nevertheless, they still struggle to capture spatial alignment and shape consistency in a unified manner. More advanced designs such as SIoU Gevorgyan (2022) integrate angular, distance, and shape components, but typically weight all samples uniformly; gradients then become dominated by easy cases while hard, tiny, low-SNR objects receive insufficient emphasis. Focaler–IoU Zhang and Zhang (2024) further reshapes the overlap signal by linearly remapping IoU within a fixed interval to favor certain ranges, but its emphasis is uniform and monotonic and it lacks angle-aware distance/shape coupling. Wise-IoU Tong et al. (2023) introduces a dynamic, non-monotonic focusing mechanism that suppresses low-quality examples, yet its default formulation is ill-suited to our setting: it deemphasizes geometric shape and aspect consistency that are crucial for sub-pixel localization of tiny defects, it normalizes outlier degree at the batch level—which can miscalibrate under low-SNR, highly imbalanced anchors and inadvertently down-weight informative hard-but-valid samples—and it omits explicit angle-aware distance coupling, yielding weaker guidance for axis-aligned center drift near overlap. Compounding these issues, conventional IoU-based objectives offer limited discrimination in the high-IoU regime where fine-grained localization matters most, leading to slow convergence and reduced accuracy.
To address these shortcomings, we propose Focaler–Wise–SIoU, a unified regression objective that couples normalized IoU scaling, comprehensive geometric penalties, and difficulty-aware modulation. The goal is to sustain strong, stable gradients across the entire overlap spectrum while explicitly promoting spatial alignment and shape fidelity for tiny objects. Concretely, we first introduce a Focaler-IoU overlap surrogate that clips and linearly rescales IoU to sharpen gradients near perfect alignment, while preserving the raw IoU for principled reweighting. We then design SIoU-style geometric penalties that combine an angle-aware, axis-normalized center-distance term with a smooth, saturating shape discrepancy, encouraging accurate alignment and aspect consistency under tiny-object noise. Finally, we incorporate a non-monotonic focal modulation (Wise) driven by an EMA-normalized IoU gap, which highlights informative, moderately hard samples and down-weights both trivial cases and extreme outliers. Raw IoU is used exclusively for difficulty estimation to retain geometric interpretability, and the overall objective remains closed-form, differentiable, and implementation-friendly. These components jointly accelerate convergence and substantially improve localization accuracy on small, low-SNR defects.
The following subsections introduce notation, define each component with its rationale, and present the final loss composition.
3.5.1 Preliminaries and Notation
Let the predicted and target bounding boxes be and , parameterized by center coordinates and width/height . We write the top-left / bottom-right corners as
The intersection size is
with area . The union area , and the standard IoU is
We also denote the center displacement by , and the side lengths of the minimum enclosing rectangle of and by
Numerical stability. We use small constants where needed: in Eq. (7) and in Sec. 3.5.6 to avoid singularities; the EMA momentum balances variance and adaptability across training phases.
3.5.2 Focaler-IoU Overlap Term
In transmission-line defect detection most objects are tiny; training benefits from non-saturating gradients even at high overlaps. To strengthen gradient sensitivity near perfect alignment while keeping the geometric meaning of Eq. (3) for reweighting, we use a clipped and linearly rescaled proxy only inside the overlap penalty:
and define
High-IoU sensitivity. Using solely in preserves gradient magnitude as while the raw IoU (Eq. (3)) is retained for difficulty modulation, maintaining geometric interpretability.
3.5.3 Angle-Weighted Distance Penalty (SIoU-style)
Let be the acute angle between and its nearest coordinate axis:
where prevents division by zero when . The angle cost in SIoU can be written as . Using and , we obtain the exact equivalence
We then define
so that the distance term below naturally takes the canonical form : larger misalignment (large ) increases the penalty before purely shrinking the radial offset.
We penalize center offsets per axis with squared, scale-normalized terms:
Because by Eq. (9), each exponential factor decreases as the (normalized) offset grows, hence is monotone in misalignment while remaining smooth around zero. Axis-wise normalization ensures scale invariance across object sizes, contributing to the overall scale invariance of the loss.
3.5.4 Shape Consistency Penalty
Tiny defects often suffer from annotation noise; a saturating yet smooth transform mitigates the dominance of extremely large aspect errors while keeping sensitivity to moderate mismatches. We adopt
For small relative deviations, is near-linear; for large deviations it saturates towards 1, and the exponent further curbs outliers while preserving gradients for mid-range shape errors. This saturating design improves robustness to annotation noise on tiny objects without letting large, noisy deviations dominate the loss.
3.5.5 Geometric Core: Overlap + Alignment + Shape
Combining the overlap penalty (Eq. (6)) with the distance and shape terms (Eqs. (10)–(11)), we obtain the SIoU-style geometric core:
The factor balances alignment/shape against overlap so that none dominates the gradient field in typical training regimes. Note again that only uses the rescaled ; all reweighting below relies on the raw IoU.
3.5.6 Wise Difficulty Modulation via EMA Normalization
We quantify sample difficulty with the normalized IoU gap
where the expectation is tracked by an exponential moving average (EMA). We use , which is stable yet responsive for shifts across training phases; in practice we add a tiny to the denominator for numerical stability. Crucially, Eq. (13) operates on the raw IoU from Eq. (3) to preserve geometric interpretability.
To emphasize moderately hard examples while avoiding overfitting to extreme outliers, we employ the following modulation:
This non-monotonic weighting targets informative moderately hard samples around , gently down-weights trivially easy cases () and prevents domination by extreme outliers as .
3.5.7 Final Objective and Batch-wise Computation
Multiplying the geometric core (Eq. (12)) by the difficulty weight (Eq. (14)) yields the Focaler–Wise–SIoU objective:
We now detail a concise batch-wise procedure that operationalizes Eqs. (5)–(11) and (13)–(14) into a practical training routine:
4 Experiments and Results Analysis
4.1 Datasets
@@
In this paper, we use the CSG-ADCD dataset. The CSG-ADCD (China Southern Grid Aerial Defective Component Dataset) is a large-scale, high-resolution UAV-based image dataset specifically developed for the detection of defects in power transmission line components, including insulators, tie wires, and poles. Constructed under authentic inspection scenarios by China Southern Power Grid, the dataset comprises 10,000 aerial images. Each image was manually annotated by a single domain expert to ensure both consistency and high-quality labeling.
@@
In summary, the CSG-ADCD dataset demonstrates strong representativeness by simultaneously addressing two critical aspects. On the one hand, it captures a comprehensive range of component defects and status indicators that are directly relevant to transmission safety, while also reflecting the complexity of real-world inspection conditions such as occlusion, cluttered backgrounds, and varying illumination. On the other hand, its pronounced dominance of small-scale objects—over 94% of all annotated instances—together with the presence of subtle and fine-grained anomalies, makes it a highly challenging and domain-specific benchmark. In this regard, CSG-ADCD not only advances research in defect detection for power systems but also provides a valuable testbed for generic small-object detection in aerial remote sensing, offering greater domain relevance than existing datasets such as VisDrone(Du et al., 2019).
@@ 4.2 Model Training and Evaluation Metrics
We implement the proposed model on the PyTorch framework and build it upon the Ultralytics RT-DETR backbone initialized with pretrained weights. Training is conducted on a workstation with an NVIDIA GeForce RTX 3090 (24 GB) running Ubuntu 20.04, Python 3.10.14, PyTorch 2.3.1, and CUDA 12.1. The model is trained for 100 epochs with a batch size of 4 and an input resolution of , using AdamW with an initial learning rate , a final learning-rate factor , momentum 0.9, and weight decay . We apply a 2000-iteration warm-up with momentum set to 0.8 and a bias learning rate of 0.1.
Validation follows the COCO protocol on the validation split: we report (AP at IoU ) and (AP averaged over IoU thresholds from 0.50 to 0.95 in steps of 0.05). Inference uses class-agnostic non-maximum suppression with an IoU threshold of and at most 300 detections per image; this IoU is unrelated to the evaluation thresholds.
@@ 4.4 Ablation Study of the Proposed Method
To analyze the impact of each component in CSG-ADCD, we conduct ablation studies by selectively enabling and disabling four key modules: EER (Edge-Enhanced ResNet), SPD (Stride-free Space-to-Depth), CSDMAM (Cross-Stage Dual-Domain Multi-Scale Attention Module), and FWS Loss (Focaler-Wise-SIoU Loss). The results in Table 2 quantify each module’s contribution across multiple evaluation metrics. Here, denotes an enabled module, and denotes a disabled one.
@@
Edge-Enhanced ResNet (EER) EER strengthens boundary-aware features and reduces background leakage. Relative to the bare baseline, precision increases from 0.369 to 0.444, recall rises from 0.177 to 0.203, and improves from 0.163 to 0.189 at an unchanged 57.0 GFLOPs. These shifts indicate fewer false positives at a fixed confidence threshold and more true positives on low-contrast contours. Importantly for the perceived instability, EER conditions the feature space so that subsequent attention behaves reliably. When EER is combined with SPD and CSDMAM, recall increases further to 0.233 and for small objects increases from 0.083 with SPD alone to 0.119 with all three. This pattern shows that EER anchors CSDMAM to genuine defect boundaries, turning a potentially volatile attention stage into a consistent contributor to small-object localization.
Space-to-Depth Convolution (SPD) Applied alone, SPD preserves fine spatial evidence while avoiding aliasing, which suppresses spurious activations. Precision increases from 0.369 to 0.486, recall holds near 0.195, and stays around 0.190, with GFLOPs rising modestly from 57.0 to 59.9. When SPD is paired with EER but without attention, recall increases from 0.195 to 0.208 and nudges from 0.190 to 0.192, while precision softens from 0.486 to 0.417, which is expected because stronger edge cues surface more borderline candidates before a selective mechanism filters them. Once CSDMAM is added, those candidates are better gated: precision recovers from 0.417 to 0.485 and recall increases from 0.208 to 0.233. Thus, SPD and EER jointly deliver detail-rich, edge-faithful inputs that convert CSDMAM from a possible source of volatility into a consistent enhancer of small-object detection, as reflected by increasing from 0.083 to 0.119 when all three are enabled.
Cross-Stage Dual-Domain Multi-Scale Attention Module (CSDMAM) The single-module ablation clarifies why a reviewer might perceive instability if attention is used in isolation. Enabling CSDMAM alone increases compute from 57.0 to 72.8 GFLOPs yet does not improve end-to-end accuracy: precision moves from 0.369 to 0.458, recall decreases from 0.177 to 0.166, and remains near baseline at 0.165. This outcome indicates that, without edge priors or detail-preserving downsampling, attention can over-weight textured background regions and depress recall. Pairwise ablations corroborate this dependency: adding CSDMAM to SPD reduces precision from 0.486 to 0.385, and adding it to EER lowers precision from 0.444 to 0.425. In contrast, when CSDMAM is combined with both EER and SPD, its intended role materializes: increases from 0.190 with SPD alone to 0.237 with all three modules, and for small objects increases from 0.083 to 0.119. Hence, CSDMAM is not intrinsically unstable; it becomes reliably beneficial once the backbone provides clean edges and SPD preserves the fine-scale evidence that attention needs to gate.
Focaler-Wise-SIoU Loss (FWS Loss) FWS Loss converts representational gains into better calibration and tighter alignment. When applied to the EER–SPD–CSDMAM stack, precision increases from 0.485 to 0.534, recall increases from 0.233 to 0.263, and increases from 0.237 to 0.275. These simultaneous gains show improved confidence shaping and stricter localization. Two interactions clarify the trade-offs. First, relative to the bare baseline, for small objects increases from 0.071 to 0.106 under the full configuration, confirming stronger tiny-target localization on average. Second, relative to the three-module model without the loss, decreases from 0.119 to 0.106 even as global metrics improve. This pattern reflects FWS Loss emphasizing medium-difficulty samples and penalizing loose boxes, which can prune marginal small-object detections at the chosen thresholds. In practice, this behavior is tunable by slightly relaxing the classification prior for the smallest anchors or moderating the modulation strength in the loss. Overall, FWS Loss interacts constructively with EER, SPD, and CSDMAM—stabilizing predictions, curbing over-confident false positives, and reinforcing the complementary effects of the modules—thereby addressing the instability observation within a co-designed pipeline.